OpenStudy (anonymous):

Thirty-six people in student council are running for the offices of president and vice-president. In how many different ways can those offices be assigned? 1,260 578 71

OpenStudy (solomonzelman):

36 P 2

OpenStudy (vishweshshrimali5):

\[\large{^{36}C_{2} \times 2!}\]

OpenStudy (vishweshshrimali5):

C - combination P - permutation

OpenStudy (anonymous):

ok so how do i do a combo

OpenStudy (vishweshshrimali5):

\[\large{^n C_{r} = \cfrac{n!}{r!(n-r)!}}\]

OpenStudy (vishweshshrimali5):

\[\large{^n P_r = \cfrac{n!}{(n-r)!}}\]

OpenStudy (vishweshshrimali5):

Actually: \[\large{^{36} P_{2} = ^{36}C_{2} \times 2!}\]

OpenStudy (anonymous):

ok thank you

OpenStudy (vishweshshrimali5):

:) Your welcome

OpenStudy (anonymous):

hmm

OpenStudy (vishweshshrimali5):

C - choosing P - arranging (another way of understanding) You can either directly arrange 2 people out of 36 - 36P2 OR Choose 2 out of 36 and then arrange them - 36C2 * 2!

OpenStudy (anonymous):

so would it be 578?

OpenStudy (anonymous):

36 choices for president one that is chosen 35 for veep by the counting principle the number of ways to do this is \(36\times 35\) do not get married to any formulas

OpenStudy (vishweshshrimali5):

1260

OpenStudy (vishweshshrimali5):

Both are same @satellite73

OpenStudy (anonymous):

all these formulas are based on formalizing the counting principle for different scenarios think don't be a slave to the formulas

OpenStudy (anonymous):

Ok, that makes a bit more scense

OpenStudy (vishweshshrimali5):

\[\large{^{36} C_{2} \times 2! = \cfrac{36!}{34!} = 36*35 = 1260}\]

OpenStudy (vishweshshrimali5):

I agree with @satellite73 . Don't use formulas unless necessary. When you can solve the problems without them then go for it

OpenStudy (anonymous):

yes, i know both are the same but it is clear that 36 choices followed by 36 choices gives \(36\times 35\) choices based on nothing more than common sense and the counting principle no formulas needed no factorials nothing of the kind

OpenStudy (anonymous):

typo there but you get the idea

OpenStudy (vishweshshrimali5):

Correct ^^^ Special cases/formulas are useless if you have the basic concepts correct

OpenStudy (vishweshshrimali5):

Though remember this: @satellite73 used `36 * 35` because it is an AND operation. You have to choose president AND vice- president

OpenStudy (solomonzelman):

vishweshshrimali5 , why do write \(\normalsize\color{blue}{ ~^{\LARGE36}P_{\LARGE2} }\) instead of \(\normalsize\color{blue}{ ~_{\LARGE36}P_{\LARGE2} }\) ?

OpenStudy (vishweshshrimali5):

Similarly, I too used `36C2 * 2!` because I have to choose 2 people out of 36 AND then I have to arrange them

OpenStudy (vishweshshrimali5):

That is notation I have always read and studied :)

OpenStudy (anonymous):

further, even \(\binom{n}{k}\) should not be formula driven imagine computing \[\binom{100}{2}\] when all that is required is \[\frac{100\times 99}{2}\]

OpenStudy (vishweshshrimali5):

Yep :)

OpenStudy (anonymous):

thanks you guys, you both have been a great help

OpenStudy (vishweshshrimali5):

Things are easier to understand and solve when you know the "WHY"

OpenStudy (vishweshshrimali5):

Your welcome @CLoudrunner :)

OpenStudy (anonymous):

@SolomonZelman i have never seen \(_nC_k\) or \(^nC_k\) in a text only \(\binom{n}{k}\)

OpenStudy (vishweshshrimali5):

I have seen both \(\large{^nC_k}\) and \(\large{\binom{n}{k}}\) - this one in foreign writers' books